OpenStudy (anonymous):

What is the diameter of a circle whose circumference is 89.49 centimeters? Use 3.14 for π. 14.25 cm 5.34 cm 29.83 cm 28.5 cm

OpenStudy (anonymous):

@___cccaroline @abb0t

OpenStudy (bradely):

pid =89.49 d =89.49/pi =28.5 cm Option D is correct Source: http://www.mathskey.com/question2answer/

OpenStudy (anonymous):

correct

OpenStudy (imstuck):

C = pi (d) 89.49 = 3.14(d) d=28.5

OpenStudy (anonymous):

Formula is 2*pi*r what is diameter in terms of r ? Try to make an equation with what is given